# Safeguarding Vision-Language Models Against Patched Visual Prompt Injectors

### v1

@@ 3.4 SmoothVLM Design
In this part, we formally introduce our design of SmoothVLM. Similar to SmoothLLM and other randomized smoothing-based methods, there are two key components: (1) distribution procedure, in which $N$ copies of the input image with random masking are distributed to VLM agents for parallel computing, and (2) aggregation procedure, in which the responses corresponding to each of the perturbed copies are collected, as depicted in Figure 1.
@@ 4.3 Efficiency
We compared the efficiency of the attack method (JIP) with our defense strategy (SmoothVLM). The JIP method focuses on reducing the loss in the embedding space of the VLM, offering a more time-efficient approach compared to the VAE, which computes the loss over the entire VLM and is thus significantly more time-consuming. Even we select JIP method, the training of an adversarial example using JIP still requires an average of 30 minutes (openai/clip-vit-base-patch32). Conversely, our defense mechanism’s time consumption depends on the number of VLM model inferences and the binary (Yes/ No) responses from the LLM. Utilizing Vicuna-30B, which has a substantial number of parameters, our method takes less than 1 minute under $N$=10, making our most resource-intensive defense approach more than 30 times faster than the fastest attack method. According to Figure 7, when we set $N$ to 10, the ASR is under 5%, indicating that the attack method is ineffective. Consequently, our approach achieves an effective and efficient success in terms of both defense robustness and time efficiency.
@@ 4.4 Compatibility
In the former section, we primarily focus on the single-patch attack. In this section, we further demonstrate the compatibility of our method with a dual patches attack. we implement two kinds of dual patch attacks (JIP, VAE) and report the defense performance of mask, swap, and replace on 300 adversarial attack examples. As shown in In Figure 7, the experimental data indicate a consistent trend where the ASR decreases as the number of samples increases for all perturbation strategies: mask, swap, and replace, across different perturbation intensities. The observed trend is further characterized by the fact that higher perturbation percentages lead to higher ASR, underscoring the defense effect in larger perturbations.

### v2

@@ 3.4 SmoothVLM Design
We formally introduce our design of SmoothVLM. Similar to SmoothLLM and other randomized smoothing-based methods, there are two key components: (1) distribution procedure, in which $N$ copies of the input image with random masking are distributed to VLM agents for parallel computing, and (2) aggregation procedure, in which the responses corresponding to each of the perturbed copies are collected, as depicted in Figure 1.
@@ 4.3 Efficiency
Smoothing-based certified defenses unavoidably increase the inference runtime Xiang et al. (2022); Cohen et al. (2019). In this section, we demonstrate that our SmoothVLM is efficient from two perspectives. We first compare the efficiency of the attack method (JIP) with our defense strategy (SmoothVLM). The JIP method focuses on reducing the loss in the embedding space of the VLM, offering a more time-efficient approach compared to the VAE, which computes the loss over the entire VLM and is thus significantly more time-consuming. Even though we select the JIP method, the optimization of a single adversarial example using JIP still requires around 30 minutes using one A100 (openai/clip-vit-base-patch32). In contrast, with the usage of a large model, Vicuna-30B, our method takes less than 1 minute under $N=10$, making our most resource-intensive defense approach more than 30x faster than the fastest attack method. As shown in Figure 7, the ASRs are consistently under 5% with $N=10$, indicating that SmoothVLM successfully guards the models. On the other hand, our method indeed offers substantial efficiency improvements. Compared to the classic randomized smoothing method Cohen et al. (2019), which necessitates 100,000 runs for a single input, our innovative smoothing approach requires only 10-20 runs per visual prompt. This represents a 10,000-fold increase in efficiency over state-of-the-art methods. Therefore, we believe SmoothVLM effectively balances efficiency and usefulness.
@@ 4.4 Compatibility
In the former section, we primarily focus on the single-patch attack. In this section, we further demonstrate the compatibility of our method with a dual patches attack. we implement both kinds of dual patch attacks and report the defense performance of mask, swap, and replace on 300 adversarial attack examples. As shown in In Figure 7, the experimental data indicate a consistent trend where the ASR decreases as the number of samples increases for all perturbation strategies: mask, swap, and replace, across different perturbation intensities. The observed trend is further characterized by the fact that higher perturbation percentages lead to higher ASR, underscoring the defense effect in larger perturbations.